Does APPLES capital structure differ from their industry average? Why or why not? Propose a change to their existing capital structure that you believe will benefit the firm. Justify your reasons for this change. Include the new values of the capital structure ratios that result from your proposed change.
Answer #1

To understand Apple’s capital structure we can look at the amount of debt and equity that the company uses to finance its operations. As of September 28, 2019 (the company’s latest financial year end) the company had total liabilities of $248,028 (in millions) and $90,488 (in millions) as assets. As such the company’s debt to equity ratio stands at 248,028/90,488 = 2.74. This indicates that the company is highly leveraged.

Apple’s capital structure does not differ significantly from the industry average and majority of the companies in this industry make use of high amount of debt to finance their operations. The only difference is that the level of leverage at Apple is higher than the industry average.

A change that will benefit the company is to repay some of its debt using its cash balance. The company had $48,844 (in millions) of cash balance as on September 28, 2019 and the company can easily use a part of this to bring down its debt level. The benefit that will accrue to the firm by reducing debt is that their financial risk will come down significantly.

I propose that the company presently make use of 50% of their cash balance to pay off their debt. So new debt amount = 248,028 – 50% of 48,844 = $223,606. In this scenario the new debt to equity ratio will be = 223,606/90,488 = 2.47. This will reduce the financial risk associated with debt for Apple and the company, in future, should look to keep on reducing its debt till it has achieved a debt equity ratio of 1.